The possible outcomes for the returns on Stock X and the returns on the market portfolio have been estimated as follows:                                                             
                                    Scenario          Stock X Market portfolio                                
                        1          9%                   9%                              
                        2          21%                 13%                            
                        3          13%                 11%                            
Each scenario is considered to be equally likely to occur. Calculate the covariance of the returns of Stock X and the market portfolio.

Group of answer choices

0.08%

1.58%

0.57%

0.04%                                                                                                         
Answer #1

it is given equally likely it means probability of occurrence is 1/3 or 0.33333

Co-variance = P(X - X)(M – M)

Where P = probability

X = return of stock X

X' = average return of stock X

M = return of market portfolio

M' = average return of market portfolio

Market Portfolio X-X -5.333 6.667 M-M -2 Scenario Probability Stock 1 0.333333339 0.33333333 21 3 0.33333333 13 Total P(X-X

from the above Table Average return of X = 43 / 3 = 14.33%

Average return on market portfolio = 33 / 3 = 11%

Co-variance = 0.08%